# Natural Language Processing for the Legal Domain: A Survey of Tasks, Datasets, Models, and Challenges

### v1

@@ 2. Background and Foundational Concepts
In this section, we will establish an essential understanding of how NLP intersects with the legal domain. Initially, we explore the characteristics of legal documents, which are fundamental to this intersection. Legal texts are known for their complex structure, specialized words. Recognizing these attributes is essential as they significantly influence the development and application of AI technologies in legal practices.
@@ 2.1. Legal Documents
Legal documents are typically written in a descriptive language and presented in an unstructured text format. They have unique features that set them apart from other fields. These documents cover a broad array of texts essential to the functioning of the legal system. They include court filings, judgments, legislation, treaties, contracts, and legal correspondence, each serving a specific purpose and following to specific formatting and content standards that reflect legal logic and hierarchy. Legal documents are Fundamental tools for lawyers, judges, and legal scholars, facilitating case analysis, legislative review, and contract drafting. They are also essential in legal education and practice and provide the basis for legal arguments and decisions. Common examples of legal documents include case law repositories, statutory databases, and collections of legal agreements. These documents are utilized in various legal processes such as drafting legal arguments, performing legal analysis, and ensuring regulatory compliance.
@@ 2.1.1. Legal language and its characteristics
Legal language is characterized by its unique features that set it apart from everyday language, primarily due to its function in the legal system. One prominent feature is its formality, where legal language often employs a more formal vocabulary and syntax to ensure precision and avoid ambiguity (Gibbons and Turell, 2008). This formality is critical, as the precise meaning of terms can have significant legal effects. Legal texts also typically utilize passive constructions and complex sentence structures to provide detailed and comprehensive descriptions (Gibbons and Turell, 2008). These constructions help clarify responsibilities and outcomes without directly attributing actions or intentions to specific parties.
Another distinctive aspect of legal language is its reliance on specialized words and phrases. This includes terms that have specific meanings within legal contexts, archaic words that are not commonly used in everyday language, and standardized phrases that have been historically embedded in legal tradition (Gibbons and Turell, 2008). This can make legal documents less accessible to non-specialists, requiring legal professionals to interpret the content accurately.
Furthermore, legal language is heavily intertextual, meaning it frequently references other legal texts, such as statutes, regulations, and case law. This characteristic ensures that legal arguments are grounded in and supported by existing legal frameworks and previous cases. The dense use of citations and references in legal documents not only supports the arguments made but also connects the document to a broader legal discourse. Such intertextuality demands that legal professionals not only understand the texts themselves but also the broader legal context in which they operate. To illustrate the intertextuality of legal language, Figure 2 shows a sample page from the Code of Federal Regulations of the United States, extracted from the WestLaw (Westlaw, [n. d.]) website. Figure 2 displays § 40.51, Labor Certification, from 22 C.F.R. § 40.51, which is part of the Code of Federal Regulations of the United States. This section falls under Title 22, governing regulations related to foreign relations, specifically detailing the requirements and procedures for labor certification. Notably, the text includes highlighted references to other legal sources, such as INA 212(a)(5). This citation refers to the Immigration and Nationality Act, specifically section 212, which outlines various conditions for inadmissibility into the United States, under subsection (a), paragraph (5). In the “Credits” part, you can see the reference “56 FR 30422, July 2, 1991,” which points to a publication in the Federal Register. Here, “56 FR” indicates the volume number, and “30422” is the page number where the document begins. The date “July 2, 1991,” marks the publication date in the Federal Register. Additionally, a sentence in subsection (b), paragraph (1) consists of 68 words, showing the length and complexity typical of legal texts.
@@
A sample page from the Code of Federal Regulations of the United States
The specific characteristics of legal language, such as its formal vocabulary, complex syntax, and extensive use of references, present many challenges for NLP. Disambiguation titles and nested entities are other issues in legal contexts (Krasadakis et al., 2024). Disambiguation titles, such as ‘The President of USA’ requires precise identification based on contextual details like time and location. Nested entities , where titles of legislative articles refer to other laws, introduce further complexity. On the other hand, legal documents are frequently provided in non-machine-readable PDF formats, complicating data extraction and processing. These challenges highlight the need for advanced and specialized NLP solutions tailored to the legal domain.
@@ 2.2.1. Introduction to Legal NLP
Legal NLP involves the application of NLP techniques to legal texts. This field is crucial as it helps automate and enhance the analysis of complex legal documents, improving efficiency and accuracy in legal research, compliance, and decision-making processes. The foundation of NLP is text, and the legal domain primarily consists of textual data (Almuslim and Inkpen, 2022), including statutes, case law, contracts, and regulations. Given the text-intensive nature of the legal field, NLP offers significant potential to transform how legal professionals interact with and utilize this vast amount of information. By leveraging advanced algorithms and Machine Learning (ML) models, Legal NLP aims to make legal texts more accessible, interpretable, and actionable.
2.2.2. Why NLP is a game-changer for the legal section
LLMs, as a part of NLP applications, such as ChatGPT (Ouyang et al., 2022), have had a significant impact since their public debut in November 2022. The legal sector, however, has been exploring the potential of AI for a longer period, applying it practically. NLP applications in the legal field are extensive, ranging from drafting client briefs and producing complex analyses from large document sets to enabling smaller firms to compete with larger ones (McGee, [n. d.]). NLP is very important in doing thorough checks when companies merge and greatly helps in legal education and learning in fast-changing fields (McGee, [n. d.]).
A notable demonstration of the capabilities of NLP applications in the legal sector occurred when GPT-4 passed the Uniform Bar Exam, underscoring the technology’s potential (Katz et al., 2024). Lawyers and law students are keenly aware of this potential, as evidenced by a LexisNexis survey released in August 2023 (LexisNexis, [n. d.]). The survey revealed that about half of all lawyers believe that LLMs will significantly transform legal practice, with nearly all (92 percent) expecting some impact. Additionally, 77 percent of respondents believe LLMs will increase the efficiency of lawyers, paralegals, and law clerks, while 63% foresee changes in how law is taught and studied. Moving forward, we will introduce specific NLP tasks in the legal domain, exploring their applications and impacts they are expected to have on the legal profession.
2.2.3. Basic foundations of NLP
The integration of NLP in the legal domain relies on several foundational techniques that enable the effective analysis and manipulation of legal texts. These fundamental methods provide the building blocks for more complex applications, transforming unstructured legal documents into structured, actionable information. By leveraging these NLP techniques, legal professionals can enhance their efficiency and accuracy, making it easier to manage and interpret vast amounts of legal data. The following section provides essential definitions of terms related to the basic foundations of NLP:
@@ 2.2.3. Basic foundations of NLP
  • Tokenization: Tokenization is the process of breaking down a sequence of text into smaller units, typically words or sub-words, known as tokens. This is a fundamental step in NLP as it allows for the structured analysis of text. In the legal domain, tokenization helps in processing and understanding lengthy documents by segmenting them into manageable pieces.
  • Word Embeddings: Word Embeddings are continuous vector representations that encode the semantic meanings of words or tokens in a high-dimensional space. These representations allow models to convert individual tokens into a format suitable for neural network processing. During training, LMs develop embeddings that capture the relationships between words, such as synonyms.
  • Transformers: Transformers (Vaswani et al., 2017) are a neural network architecture designed to convert input sequences into output sequences by understanding the context and relationships among the elements of the sequence. For instance, given the input “What is the color of the sky?”, a transformer model internally processes and identifies connections among the words ‘color’, ‘sky’, and ’blue’. This understanding enables it to produce the output: “The sky is blue” (Amazon, [n. d.]). Transformer models enhance this process through a self-attention mechanism. This mechanism allows the model to analyze different parts of the sequence simultaneously instead of sequentially, helping it identify which parts are most significant.
  • PLMs: PLMs are trained on extensive corpora in a self-supervised manner, which involves tasks like recovering incomplete input sentences or auto-regressive language modeling. These models, such as Bidirectional Encoder Representations from Transformers (BERT) (Devlin et al., 2019) and RoBERTa (Liu et al., 2019), are initially trained on large-scale general text datasets. After pre-training, they can be fine-tuned for specific downstream tasks in the legal domain, adapting them to comprehend and process legal language for applications like document classification and information extraction.
  • Question Answering (QA): QA system is a type of NLP solution designed to answer questions posed in natural language. These systems take a user’s query and, by extracting relevant information from a dataset provide a relevant and informative response.
  • NER: NER is the task of identifying mentions of specific entities within a text that belong to predefined categories such as persons, locations, organizations, and more. NER is a fundamental component for many NLP applications, including question answering, text Summarization, and machine translation (Li et al., 2022b).
  • Information Retrieval (IR): IR involves the process of obtaining relevant information from large collections of unstructured legal texts, such as case laws, statutes, contracts, and regulations. The goal of IR is to provide users with the most relevant documents and data in response to specific queries.
  • Multi-task Learning (MTL): MTL is an approach in ML where a model is trained on multiple related tasks simultaneously or utilizes auxiliary tasks to enhance performance on a specific task. By learning from diverse tasks, MTL enables models to capture generalized and complementary knowledge, improving robustness and addressing data scarcity, particularly for low-resource tasks. MTL’s ability to share implicit knowledge across tasks often leads to performance gains and more efficient models, making it a valuable strategy for building robust and adaptable systems in NLP and other domains (Chen et al., 2024b).
@@
  • Parameter-Efficient Fine-Tuning (PEFT): PEFT is a method for adapting PLMs that involves freezing the majority of the model’s parameters and only updating a small subset. This approach significantly reduces the computational resources and time required for fine-tuning, making it particularly effective in resource-limited scenarios, while still achieving competitive performance in tasks like text generation (Li et al., 2024a).
  • Retrieval-Augmented Generation (RAG): RAG is an advanced AI framework that enhances text creation by merging traditional information retrieval systems, with the generative power of LLMs. This integration allows the AI to access additional knowledge sources while utilizing its advanced language capabilities.
@@ 2.2.4. Key Publications and Conferences in Legal NLP
Several leading journals focus on the intersection of AI, NLP, and the legal domain. “Artificial Intelligence and Law”, published by Springer, is a leading journal that features research articles on legal reasoning, legal IR, and legal knowledge representation. Additionally, the “Journal of Law and Information Technology”, published by Taylor & Francis , focuses on the application of information technology in law, including research AI.
Conferences significantly advance research and promote collaboration in Legal NLP. The International Conference on Artificial Intelligence and Law (ICAIL) is a biennial event that showcases advances in AI applications in the legal domain, including NLP and ML. The Conference on Legal Knowledge and Information Systems (JURIX) is an annual event that focuses on legal informatics and NLP technologies.
In Legal NLP, there are some good works which are presented in workshops. The Workshop on Automated Semantic Analysis of Information in Legal Texts delves into NLP and semantic analysis of legal texts. The Natural Legal Language Processing (NLLP) Workshop provides a platform for discussing NLP technologies tailored for legal texts and is often part of major NLP conferences. The EXplainable AI in Law (XAILA) Workshop focuses on the explainability of AI systems in legal contexts, aiming to improve transparency and trust in AI applications. The Competition on Legal Information Extraction/Entailment (COLIEE) is an annual event that challenges participants to develop innovative solutions for legal information extraction and entailment tasks.
@@ 3.1. Search Strategy
  • Query 2: (“Legal” AND (“Named Entity Recognition” OR “NER” OR “Document Summarization” OR “Text Classification” OR “Document Classfication” OR “Judgment Prediction” OR “Question Answering” OR “Corpus” OR “Language Model”))
Our search covered publications with the following date ranges for each specific NLP task: LQA from 2020-2024, LJP from 2017-2024, LTC from 2018-2023, LDS from 2016-2024, Legal NER from 2010-2022, legal corpora from 2021-2024, and Legal LMs from 2020-2024. This ensured that we included recent advancements. Peer-reviewed journal articles and high-quality conference proceedings were prioritized, along with a secondary consideration of non-peer-reviewed sources, such as arXiv articles, where relevant.
@@ 3.2. Study Selection
A total of 148 studies were initially identified from the database search. To refine this list, we applied a manual reviewing and. This process involved:
@@
    Title and Abstract Screening: We reviewed the titles and abstracts of all retrieved studies to assess their relevance to the predefined legal NLP tasks. Studies unrelated to the core legal NLP and its tasks were excluded at this stage.
@@
    Final Selection: From the original 148 studies, 127 were included in the final list, selected based on their direct relevance to key NLP tasks within the legal sector and their methodological quality, as well as their engagement with existing literature in the field.

### v2

@@ 3. Background and Foundational Concepts
In this section, we will establish an essential understanding of how NLP intersects with the legal domain. Initially, we explore the characteristics of legal documents, which are fundamental to this intersection. Legal texts are known for their complex structure, specialised words. Recognising these attributes is important because they influence the development and use of AI in legal practices.
@@ 3.1. Legal Documents
Legal documents are typically written in a descriptive language and presented in an unstructured text format. They have unique features that set them apart from other fields. These documents cover a broad array of texts essential to the functioning of the legal system. They include court filings, judgements, legislation, treaties, contracts, and legal correspondence, each serving a specific purpose and following to specific formatting and content standards that reflect legal logic and hierarchy. Legal documents are Fundamental tools for lawyers, judges, and legal scholars, facilitating case analysis, legislative review, and contract drafting. They are also essential in legal education and practice and provide the basis for legal arguments and decisions. Common examples of legal documents include case law repositories, statutory databases, and collections of legal agreements. These documents are utilised in various legal processes such as drafting legal arguments, performing legal analysis, and ensuring regulatory compliance.
@@ 3.1.1. Legal language and its characteristics
Legal language is characterised by its unique features that set it apart from everyday language, primarily due to its function in the legal system. One prominent feature is its formality, where legal language often employs a more formal vocabulary and syntax to ensure precision and avoid ambiguity (Gibbons and Turell, 2008). This formality is important, as the meaning of terms can have legal effects. Legal texts also typically utilise passive constructions and complex sentence structures to provide detailed and comprehensive descriptions (Gibbons and Turell, 2008). These constructions help clarify responsibilities and outcomes without directly attributing actions or intentions to specific parties.
Another distinctive aspect of legal language is its reliance on specialised words and phrases. This includes terms that have specific meanings within legal contexts, archaic words that are not commonly used in everyday language, and standardised phrases that have been historically embedded in legal tradition (Gibbons and Turell, 2008). This can make legal documents less accessible to non-specialists, requiring legal professionals to interpret the content accurately.
Furthermore, legal language is heavily intertextual, meaning it frequently references other legal texts, such as statutes, regulations, and case law. This characteristic ensures that legal arguments are grounded in and supported by existing legal frameworks and previous cases. The dense use of citations and references in legal documents not only supports the arguments made but also connects the document to a broader legal discourse. Such intertextuality demands that legal professionals not only understand the texts themselves but also the broader legal context in which they operate. To illustrate the intertextuality of legal language, Figure 1 shows a sample page from the Code of Federal Regulations of the US, extracted from the Electronic Code of Federal Regulations [^1] https://www.ecfr.gov. Figure 1 displays § 40.51, labour Certification, from 22 C.F.R. § 40.51, which is part of the Code of Federal Regulations of the US. This section falls under Title 22, governing regulations related to foreign relations, specifically detailing the requirements and procedures for labour certification. Notably, the text includes highlighted references to other legal sources, such as INA 212(a)(5). This citation refers to the Immigration and Nationality Act, specifically section 212, which outlines various conditions for inadmissibility into the US, under subsection (a), paragraph (5). In the “Credits” part, you can see the reference “56 FR 30422, July 2, 1991,” which points to a publication in the Federal Register. Here, “56 FR” indicates the volume number, and “30422” is the page number where the document begins. The date “July 2, 1991,” marks the publication date in the Federal Register.
@@
A sample page from the Code of Federal Regulations of the US
The specific characteristics of legal language, such as its formal vocabulary, complex syntax, and extensive use of references, present many challenges for NLP. Disambiguation titles and nested entities are other issues in legal contexts (Krasadakis et al., 2024). Disambiguation titles, such as ‘The President of USA’ requires precise identification based on contextual details, such as time and location. Nested entities, where titles of legislative articles refer to other laws, introduce further complexity. On the other hand, legal documents are frequently provided in non-machine-readable PDF formats, complicating data extraction and processing. Additionally, the variation in legal reasoning, which includes rule-based, analogical, and evidentiary arguments, along with changes in legal standards, creates challenges for applying conventional NLP models (Freeman Engstrom and B. Gelbach, 2021). Further elements, such as the peculiar use of punctuation affecting text segmentation and the frequent presence of digits and numbers, can also disrupt traditional NLP pipelines. These challenges show the need for specialised NLP solutions tailored to the legal domain.
3.1.2. Domains with Shared Characteristics
Other domains exhibit features similar to those of legal texts, such as specialised vocabularies, large-scale corpora, and cross-referencing. In the medical domain, texts—including clinical notes, patient records, and research articles—rely on domain-specific terminology, diagnostic labels, and biological taxonomies, and often involve case-based analyses akin to legal reasoning. Similarly, software documentation and source code contain specialised programming terms and algorithmic descriptions, and are characterised by large repositories with numerous cross-references to libraries and functions.
@@ 3.2.1. Introduction to legal NLP
Legal NLP involves the application of NLP techniques to legal texts. This field is crucial as it can help automate and enhance the analysis of complex legal documents, compliance, and decision-making processes (Huang et al., 2024). The foundation of NLP is text, and the legal domain primarily consists of textual data (Almuslim and Inkpen, 2022), including statutes, case law, contracts, and regulations. Given the text-intensive nature of the legal field, NLP offers potential to change how legal professionals interact with and use this information. By leveraging advanced algorithms and Machine Learning (ML) models, legal NLP aims to make legal texts more accessible, interpretable, and actionable (Huang et al., 2024).
3.2.2. Basic foundations and concepts of NLP
The integration of NLP in the legal domain relies on foundational techniques that enable the processing and analysis of legal texts. These techniques form the basis for various applications, transforming unstructured legal documents into structured, actionable information. This section introduces key NLP methods, machine learning paradigms, and text retrieval technique.
  1. (1)
    Core NLP methods:
@@ 3.2.2. Basic foundations and concepts of NLP
    • Tokenization: Tokenization is the process of breaking text into smaller units, typically words or sub-words, known as tokens. This is a fundamental step in NLP, allowing structured text analysis. In legal NLP, tokenization can facilitate processing lengthy documents by segmenting them into manageable parts. Tokenization prepares the input text for numerical representation through word embeddings.
    • Word Embeddings:It represent words as continuous vectors in a high-dimensional space, capturing their semantic relationships. These embeddings enable models to encode word meanings and relationships, which are essential for tasks such as legal text similarity analysis and document classification. Word embeddings provide a numerical representation of tokens, which is essential for transformer-based models to process contextual relationships.
    • Transformers: The transformer architecture is a NN model designed to process text by first identifying connections between input words using self-attention. In this process, the model computes attention weights (which determine the importance of each word based on its context) for the entire input sequence. Only after establishing these connections does it generate output representations. Transformers operate on word embeddings and refine these embeddings by incorporating context from the entire sequence.
    • PLMs: PLMs, such as Bidirectional Encoder Representations from Transformers (BERT) (Devlin et al., 2019), are transformer-based models that have been trained on large-scale text corpora in a self-supervised manner. These models learn linguistic patterns and can be fine-tuned for specific legal NLP tasks. PLMs leverage transformer architectures and pre-trained embeddings to perform NLP tasks.
  2. (2)
    ML paradigms for NLP:
    • Multi-task Learning (MTL): MTL is an approach where a model learns multiple related tasks simultaneously, leveraging shared knowledge across tasks. This technique improves model robustness and efficiency, particularly in data-scarce legal NLP applications (Chen et al., 2024b).
@@
    • Parameter-Efficient Fine-Tuning (PEFT): PEFT is a method for adapting PLMs that involves freezing the majority of the model’s parameters and only updating a small subset. This approach significantly reduces the computational resources and time required for fine-tuning, making it particularly effective in resource-limited scenarios, while still achieving competitive performance in tasks, such as text generation (Li et al., 2024a).
    These machine learning paradigms enhance the performance of core NLP methods, particularly PLMs, by tailoring models to specific tasks and optimising their training efficiency.
  3. (3)
    Text retrieval technique:
    • Retrieval-Augmented Generation (RAG): RAG combines traditional Information Retrieval (IR) methods with generative NLP models, allowing systems to retrieve external knowledge before generating responses.
    Text retrieval techniques, such as RAG complement the core NLP methods by providing additional context and external information that enhances the generation and refinement of texts.
3.2.3. LLMs as an application of NLP in the legal sector
LLMs are a category of Deep Learning (DL)-based NLP models trained on extensive text corpora to process and generate human-like language. These models, typically based on Transformer architectures, learn linguistic patterns from diverse sources, enabling them to perform tasks, such as text summarisation, document analysis, and Question Answering (QA). LLMs, as part of NLP applications, have gained widespread attention, particularly with the release of ChatGPT in November 2022 (Ouyang et al., 2022). However, the legal sector has been exploring AI-driven solutions for much longer, applying NLP techniques to automate legal processes. NLP applications in the legal field include drafting client briefs, analysing large document sets, and enabling smaller firms to compete more effectively with larger ones (McGee, [n. d.]). NLP is very important in doing thorough checks when companies merge and greatly helps in legal education and learning in fast-changing fields (McGee, [n. d.]).
A notable demonstration of NLP’s potential in the legal sector occurred when GPT‑4 was reported to have passed the Uniform Bar Exam near the 90th percentile, underscoring the technology’s promise (Katz et al., 2024).However, subsequent analyses by Martínez (2024) suggest that its actual performance may be considerably lower, possibly around the 48th percentile overall and 15th percentile on essays. Meanwhile, further research reveals that although ChatGPT can achieve moderate success on certain legal classification tasks, smaller fine-tuned models still outperform it by about 30 percentage points (Chalkidis, 2023). Another study highlights a tendency for LLMs to hallucinate legal content at high rates—up to 58% in some cases—raising reliability concerns for complex tasks (Dahl et al., 2024). Despite these drawbacks, lawyers and law students remain conscious of the broader impact of such models: a recent LexisNexis survey (LexisNexis, [n. d.]) shows that about half of all lawyers believe LLMs will transform legal practice, with 92% anticipating some impact. Furthermore, 77% of respondents foresee efficiency gains for legal professionals, and 63% predict changes in how law is taught and studied.
@@ 3.2.4. Key publications and conferences in legal NLP
Several leading journals focus on the intersection of AI, NLP, and the legal domain. “Artificial Intelligence and Law”, published by Springer, is a leading journal that features research articles on legal reasoning, legal IR, and legal knowledge representation. A recent special issue, Applications and Evaluation of LLMs in the Legal Domain, examines the use of LLMs in legal tasks such as summarisation, judgement prediction, and contract drafting, while also addressing concerns, such as bias, misinformation, and regulatory compliance. Additionally, the “Journal of Law and Information Technology” focuses on the application of information technology in law, including research AI.
Conferences significantly advance research and promote collaboration in legal NLP. The International Conference on Artificial Intelligence and Law (ICAIL) is a biennial event that showcases advances in AI applications in the legal domain, including NLP and ML. The Conference on Legal Knowledge and Information Systems (JURIX) is an annual event that focuses on legal informatics and NLP technologies. In addition to these dedicated venues, prominent legal NLP research has also been published in broader AI, NLP, and IR conferences, including NeurIPS, ACL (and its associated workshops), NAACL, EMNLP, SIGIR, IJCAI, AAAI, and LREC.
In legal NLP, there are some good works which are presented in workshops. The workshop on Automated Semantic Analysis of Information in legal texts focuses on NLP and semantic analysis of legal texts. The international workshop on Juris-Informatics (JURISIN) brings together researchers from law, social science, and technology to discuss foundational and practical issues at the intersection of legal theory and informatics. The Natural Legal Language Processing (NLLP) workshop provides a platform for discussing NLP technologies tailored for legal texts and is often part of major NLP conferences. The EXplainable AI in Law (XAILA) workshop focuses on the explainability of AI systems in legal contexts, aiming to improve transparency and trust in AI applications. The Competition on Legal Information Extraction/Entailment (COLIEE) is an annual event that challenges participants to develop innovative solutions for legal information extraction and entailment tasks. Additionally, the Legal Track at the TREC, which ran from 2006 to 2011, focused on evaluating IR techniques for legal document review, providing a venue for researchers to test and compare methods in the context of e-discovery. Its contributions include benchmark datasets and evaluation metrics that were used to assess retrieval performance in legal text processing.
@@ 4.1. Search Strategy
  • Query 2: (“Legal” AND (“Named Entity Recognition” OR “NER” OR “Document Summarisation” OR “Text Classification” OR “Document Classification” OR “Judgement Prediction” OR “Question Answering” OR “Corpus” OR “Language Model” OR “Argument Mining”))
Our search covered publications with the following date ranges for each specific NLP task: LQA from 2020-2024, LJP from 2017-2024, LTC from 2018-2023, LDS from 2016-2024, legal NER from 2010-2022, LAM from 2009-2024, legal corpora from 2021-2024, and legal LMs from 2020-2024. This ensured that we included recent advancements. Peer-reviewed journal articles and high-quality conference proceedings were prioritised, along with a secondary consideration of non-peer-reviewed sources, such as arXiv articles, where relevant.
@@ 4.2. Study Selection
A total of 154 studies were initially identified from the database search. To refine this list, we applied a manual reviewing and. This process involved:
@@
    Title and Abstract Screening: We reviewed the titles and abstracts of all retrieved studies to assess their relevance to the predefined legal NLP tasks. Studies unrelated to the core legal NLP and its tasks were excluded.
@@
    Final Selection: From the original 154 studies, 133 were included in the final list, selected based on their direct relevance to key NLP tasks within the legal sector and their methodological quality, as well as their engagement with existing literature in the field.